OpenStudy (anonymous):

solve the equation. check the solution.-4/x+1=-1/x+5

OpenStudy (eric_d):



OpenStudy (eric_d):

Cross multiplication

OpenStudy (anonymous):

\[\frac{ -4 }{ x+1 } = \frac{ -1 }{ x+5}\] multiplying the both sides by their respective denominators : \[-4x -20 = -x -1\] which isn't pssible therefore: \[\frac{ -4 }{ x+1 } \neq \frac{ -1 }{ x+5 }\]

OpenStudy (anonymous):

@That_Pro_Angler

OpenStudy (anonymous):

so then what would my answer be? @Adjax

OpenStudy (anonymous):

your answer is that : \[\frac{ -4 }{ x+1 } \neq \frac{ -1 }{ x+ 5 }\] I already gave the working above

OpenStudy (anonymous):

but thats not one of the answer choices :(

OpenStudy (anonymous):

so what are the choices then?

OpenStudy (anonymous):

A -19/4 b 1/3 c -19/3 d 2

OpenStudy (anonymous):

ok I got it..what you meant : \[-4x-20 = -x-1 => -3x =19=> x = \frac{ 19 }{ -3}\]

OpenStudy (anonymous):

Oh okay i see. Thanks :)